OpenStudy (anonymous):

A solution is made by dissolving 21.5 grams of glucose (C6H12O6) in 255 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m? Show all of the work needed to solve this problem.

OpenStudy (anonymous):

I would like to understand how to do this fully please...

OpenStudy (wolfe8):

I'm guessing the freezing constant with unit C/m means celsius per molal?

OpenStudy (wolfe8):

I'm not sure but if so, would you just multiply the molality with that constant? Might wanna wait for someone else. But welcome to OpenStudy

OpenStudy (anonymous):

I would divide the grams of glucose by the grams of water then multiply it by the constant. Im not too sure, correct me if im wrong.

OpenStudy (anonymous):

Im not sure either..... would it look like this 21.5/255 x -1.8

OpenStudy (anonymous):

yes, i would think so, but dont take my word though, im not sure....

OpenStudy (anonymous):

okaaay. well thank u anyway :)

OpenStudy (wolfe8):

Well if I'm right, you find the volume of water using its density. Then divide the number of moles of glucose by the volume of water. Then multiply by the constant. That is, if I understand the constant correctly.

OpenStudy (anonymous):

Try this @dwright25 I thought it is a good source for help

OpenStudy (wolfe8):

Oh you need more with that equation. Refer to the Calculation section http://en.wikipedia.org/wiki/Freezing-point_depression

OpenStudy (anonymous):

thanks u guys

OpenStudy (anonymous):

No problem, always available for help

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{\theta}\) where, \(\theta=b\) when finding BP, and \(\theta=f\) when finding freezing point (By that i mean use the appropriate constant.) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362